OpenStudy (raffle_snaffle):

How long does it take to triple an investment of $23360 if the investment pays only simple interest at the rate of 5.18% per annum?

OpenStudy (raffle_snaffle):

wait...

OpenStudy (raffle_snaffle):

so our principal is 23360?

OpenStudy (raffle_snaffle):

@mathmale

OpenStudy (raffle_snaffle):

Where do I start? Ask me questions that lead me to the solution.

OpenStudy (retireed):

Use this equation.... Future Value = Present Value ( 1 + Interest Rate ) ^ N you know everything except N, so solve for N. Is this clear?

OpenStudy (raffle_snaffle):

Yes, I've been using that equation for this hw. However, I am having difficulty determining what P and F are. What is principal? Money you start with? The money you barrow or loan?

OpenStudy (retireed):

Those are good questions. You want a investment of $23360 to triple in value, so ideally you know the PV is 23360 and the FV is three times that or 70080. However it is just FV/PV=3 so the equation becomes.... 3= (1+IR)^N

OpenStudy (raffle_snaffle):

I had it setup like this 70080 = 23360(0.0518)*n solve for n Why is it FV/PV? Is it a ratio?

OpenStudy (raffle_snaffle):

oh shoot that is the same equation.

OpenStudy (raffle_snaffle):

oh wait I didn't use the compound interest formula...

OpenStudy (retireed):

To solve your equation divide both sides by 23360. Also you forgot the 1 in (1+0.0518),

OpenStudy (raffle_snaffle):

Well I thought we were looking at simple interest that is why I used F = P*i*n

OpenStudy (raffle_snaffle):

because it says simple interest

OpenStudy (retireed):

Oh, I think you are right. It does say simple interest rate. I was thinking it was simple because it wasn't compounded monthly.

OpenStudy (raffle_snaffle):

okay cool

OpenStudy (raffle_snaffle):

I got n = 57.91 years

OpenStudy (raffle_snaffle):

so about 58 years if we round

OpenStudy (retireed):

but it says "simple interest at the rate of 5.18% per annum" so is the interest calculated per annum and added in or not. Did they given an answer for this question? I got the 57.91 year simple interest and around 21.74 years compounded yearly.

OpenStudy (raffle_snaffle):

no, I can find out now. I only get three attempts for each hw problem. Let me check because this was my last problem.

OpenStudy (raffle_snaffle):

nope not correct

OpenStudy (raffle_snaffle):

Well the question changed.... How long does it take to quadruple an initial investment of $40738 if the investment pays only simple interest at the rate of 10.97% per annum?

OpenStudy (retireed):

Try this and see if you get the right answer... log(4) / log (1.1097) = 13.318 years How accurate do they want you to be? How many decimal points?

OpenStudy (retireed):

If the answer is right I can explain it to you.

OpenStudy (raffle_snaffle):

Is this to the original question or the current question. After I attempt the hw once all the questions change. So I need help with this one How long does it take to quadruple an initial investment of $40738 if the investment pays only simple interest at the rate of 10.97% per annum?

OpenStudy (retireed):

My previous answer is for the new question. 13.318 years I think they really wan the interest compounded yearly.

OpenStudy (raffle_snaffle):

Okay. So how about the question I posted just recently?

OpenStudy (retireed):

YES ----- >>>> 13.318

OpenStudy (raffle_snaffle):

lol sorry. Let me do the work

OpenStudy (raffle_snaffle):

this is compound or simple?

OpenStudy (raffle_snaffle):

I got it... It was compound

OpenStudy (raffle_snaffle):

can you help me with another problem plz?

OpenStudy (retireed):

Yes that is what I am trying to figure out. Enter my answer 13.3 years, see if it is correct, then I will know what they are asking for.

OpenStudy (raffle_snaffle):

I can't... Lol I have to answer all 7 questions before I can submit. I only get 3 times to submit so I have to redo 6 more problems....

OpenStudy (raffle_snaffle):

here is my other question... For the cash flow diagram below, which equation demonstrates the correct functional notation that will solve for the unknown Z? Z = -5000(F/P,8%,4) - 3500(F/P,8%,1) + 2000(F/P,8%,2) Z = -5000 + 2000 - 3500(F/P,8%,1) Z = -5000 + 2000(F/P,8%,2) -3500(F/P,8%,1) Z = -5000 + 2000(P/F,8%,2) - 3500(P/F,8%,3) Z = -8500(F/P,8%,4) + 2000(F/P,8%,2)

OpenStudy (retireed):

Okay much easier. What do you think?

OpenStudy (raffle_snaffle):

let me stair at it more

OpenStudy (retireed):

Very good take you time.

OpenStudy (raffle_snaffle):

does the interest rate account for all the money borrowed or lend?

OpenStudy (raffle_snaffle):

does it account for the 5 years is what I am asking

OpenStudy (retireed):

Yes. In the real world, not many people will lend money without charging you interest.

OpenStudy (raffle_snaffle):

a?

OpenStudy (retireed):

It is A, very good.

OpenStudy (raffle_snaffle):

wait a min... why is some of them say P/F and F/P

OpenStudy (raffle_snaffle):

should it be P/F

OpenStudy (raffle_snaffle):

nvm that is notation. i need to read more about that...

OpenStudy (retireed):

??? I'm not seeing any question.

OpenStudy (raffle_snaffle):

Jana just received a graduation gift of $43385 from her parents. She decides to deposit it into an account that pays a guaranteed 4% per year, compounded annually. How much will she have in the account when she retires in 28 years?

OpenStudy (raffle_snaffle):

Here is my next question? Do you want me to make a new thread so I can give you a medal?

OpenStudy (raffle_snaffle):

Actually I know how to solve this... we use F = P(1+i)^n

OpenStudy (retireed):

Metal is not important, but thanks for asking. This one uses the equation from before.... Future Value = Present Value ( 1 + Interest Rate ) ^ N You know everything, so calculate the future value.

OpenStudy (raffle_snaffle):

F = P(1+i)^n = 43385(1+0.04)^28 = 130098.74

OpenStudy (retireed):

Your typing beat mine, good job.

OpenStudy (raffle_snaffle):

Davis is saving money for a bicycle trip around the world. He would like to have $23651 saved in 5 years. To reach his goal, how much does Davis need to deposit today in an account paying 7% per year compounded annually?

OpenStudy (raffle_snaffle):

Next question^^^

OpenStudy (retireed):

Your previous answer was right. This question is very similar Future Value = Present Value ( 1 + Interest Rate ) ^ N now solve for the present value.

OpenStudy (raffle_snaffle):

so future value is 23651?

OpenStudy (retireed):

yes

OpenStudy (raffle_snaffle):

P = 16862.84

OpenStudy (retireed):

correct

OpenStudy (raffle_snaffle):

For his trip, Davis deposited $13479 into an account paying 10% per year, compounded annually. Unfortunately, he had a medical emergency at the end of year 3 and had to withdraw $2114. At the end of year 7 he received a small inheritance and deposited $3408 into his account. How much did Davis have in his account at the end of 12 years?

OpenStudy (raffle_snaffle):

I think I know how to solve this one. Let me do the work on paper and attach my answer.

OpenStudy (retireed):

Yes this was is easy, but a lot more messy ...

OpenStudy (raffle_snaffle):

F=13479(1+0.10)^3 = 17940.549

OpenStudy (raffle_snaffle):

17940.549 - 2114 = 15826.549

OpenStudy (retireed):

Yes. Yes.

OpenStudy (retireed):

It is very good doing in parts and checking. I already have the final answer, but it seems too high to be correct. Keep going with what you are doing.

OpenStudy (raffle_snaffle):

F = 15826.549(1+0.10)^2 = 19150.12439

OpenStudy (raffle_snaffle):

P = 19150.12439 + 3408 = 22558.12429

OpenStudy (raffle_snaffle):

F=22560.07

OpenStudy (retireed):

Wait a second let me check something

OpenStudy (retireed):

Why did you use TWO years

OpenStudy (retireed):

F = 15826.549(1+0.10)^2 = 19150.12439 Why two years? The end of year 7 from year 3, it should be FOUR years. Right?

OpenStudy (raffle_snaffle):

oops....

OpenStudy (raffle_snaffle):

F = 15826.549(1+0.10)^4 = 23171.65 P = 23171.65 + 3408 = 26579.65 F = 26579.65(1+0.10)^7 = 51796.22

OpenStudy (retireed):

Almost let check....

OpenStudy (raffle_snaffle):

That is not the answer????

OpenStudy (retireed):

F = 26579.65(1+0.10)^7 = 51796.22 This is wrong number of years it is not 7 what should it be?

OpenStudy (raffle_snaffle):

5

OpenStudy (retireed):

End of year 7, he made the deposit. How much did he have at the end of year 12? Yes 5.

OpenStudy (raffle_snaffle):

F = 42806.79

OpenStudy (retireed):

Yep, that' what I got. You have got to be very careful with the YEARS. Other than that you did great.

OpenStudy (raffle_snaffle):

Kreskin borrowed $1500 at an annual interest rate of 10% and paid off the loan (principal and interest), after several years with a $1997 check. How many years did it take Kreskin to repay the loan?

OpenStudy (raffle_snaffle):

Okay I will keep that in mind.

OpenStudy (retireed):

You got this one....

OpenStudy (raffle_snaffle):

1997 = 1500(1+0.10)^n solve for n

OpenStudy (raffle_snaffle):

3 years

OpenStudy (retireed):

Yep

OpenStudy (raffle_snaffle):

If a sum is worth $615 today and it will be worth $731 two years from now, how much was it worth one year ago?

OpenStudy (raffle_snaffle):

(615+731)/2 = 673

OpenStudy (raffle_snaffle):

673-615 = 58

OpenStudy (retireed):

I don't know but I don't think it is that simple.

OpenStudy (retireed):

You need to calculate the interest rate some how.

OpenStudy (raffle_snaffle):

615 - 58 = 557

OpenStudy (retireed):

FV = PV (1+IR)^N you know everything but IR

OpenStudy (raffle_snaffle):

673 for every 1 year

OpenStudy (retireed):

Once you know the IR you can find the value at end of year one.

OpenStudy (retireed):

Don't forget the first year will earn less than the second year. Compound interest is still in affect.